OpenStudy (anonymous):

multiply. simplify the product 2x/x+4 times x-1//3

OpenStudy (anonymous):

some typo here yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what does double slash mean

OpenStudy (anonymous):

they r fractions

OpenStudy (anonymous):

double slash means typing mistake for single slash.

OpenStudy (anonymous):

\[\frac{2x}{x+4}\times \frac{x-1}{3}\]?

OpenStudy (anonymous):

2x over x+4 times x-1 over 3 sorry about the double slash trying to type to fast

OpenStudy (anonymous):

ok just multiply using the distributive law \[\frac{2x\times(x-1)}{(x+4)\times3}\]

OpenStudy (anonymous):

\[ \frac{2x^2-2x}{3x+12}\]